OpenStudy (anonymous):

Factor the trinomial. 6y^2 + 7y + 2

OpenStudy (anonymous):

6y^2 + 7y + 2 6y^2 + 3y + 4y +2 3y(2y+1) + 2(2y+1) Answer: 3y +2= 2y+1= -2/3 -1/2